# On the influence of electrolytic gradient orientation on phoretic transport in dead-end pores

### v1

@@ Abstract
Electrolytic diffusiophoresis refers to directional migration of colloids due to interfacial forces that develop in response to local electrolytic concentration ($c$) gradients. This physicochemical transport provides an efficient alternative in numerous microscale applications where advection-induced transport is infeasible. Phoretic withdrawal and injection in dead-end pores can be controlled by orienting salt gradients into or out of the pore; however, the extent to which this orientation influences spatiotemporal transport patterns is not thoroughly explored. In this study, we find that it has a significant influence: colloidal withdrawal in solute-out mode ($\beta=c_{\infty}/c_{\text{pore}}<1$) is faster and shallower, whereas the solute-in mode enables deeper withdrawal. Similarly, solute-out injection features rapidly propagating wavefronts, whereas the solute-in mode ($\beta>1$) promotes uniform and gradual injection. Each mode’s transport is found to evolve and persist over different time scales. We characterize the performance of these modes and find that while persistence of the solute-out mode strengthens with a growing electrolytic gradient [$\sim\ln(\beta^{-0.4})$], solute-in mode diminishes and eventually its persistence is insensitive to $\beta$. We also incorporate the variable mobility model to examine the impact of large zeta potentials, which intensifies the transport of solute-out mode further and weakens the solute-in mode. Additionally, we investigate how osmotic flows of the two modes affect injection and withdrawal patterns. We find that osmosis-induced mixing can counterintuitively inhibit injection effectiveness in solute-out mode. These insights bring attention to the distinctions between different phoretic transport modes and contribute to the rational design and setup of electrolytic gradients in numerous microscale applications.
I Introduction
Microscopic colloids in a molecular bath experience potential forces from intermolecular interactions within the Debye screening region near their surface. External concentration gradients disrupt the equilibrium of thermal and interactive forces, inducing a diffusioosmotic (DO) slip velocity. This slip on a freely suspended colloid drives a diffusiophoretic (DP) motion, where it migrates parallel to chemical gradients $\propto\nabla c$ Derjaguin et al. [1947], Anderson [1989]. When the solute is an electrolyte, differential ion diffusion creates electrostatic gradients ($\nabla\psi$) alongside chemical gradients. This generates a local electric field that modifies the osmotic/phoretic contribution, proportional to $\nabla\log c$ Prieve et al. [1984], Velegol et al. [2016], Ault and Shin [2024]. This motion is mediated by diffusive and electrostatic solute-surface interactions, encapsulated in chemo- and electro-phoretic mobility parameters.
The resurgence of these phoretic and osmotic phenomena after decades of initial work is attributed to the rise of micro- and nano-fluidic technologies, which enable precise control over fluid and particle transport Shi et al. [2016]. These physicochemical effects can also be leveraged for designing bioassays that trap vesicles Rasmussen et al. [2020] & transport DNA Katzmeier and Simmel [2024], performing low-cost zeta potentiometry Shin et al. [2017a], Ault et al. [2019], and designing hydrophobic nanopores that can act as osmotic diodes for rectified transport Picallo et al. [2013]. Furthermore, potential discoveries of osmotic and phoretic phenomena have recently emerged within applications and technologies that have existed for decades. Examples include, dirt transport in constricted fibre pores during conventional laundering Shin et al. [2018], phoresis-induced fouling, osmotic flows in water filtration membranes Sablani et al. [2001], Shin et al. [2017b], and low-salinity water flooding for tertiary oil recovery Kar et al. [2015], Velegol et al. [2016], Lager et al. [2008].
@@ 0.2-1 $\mu$m polystyrene and
Reassessed Staffeld and Quinn [1989], Ault et al. [2017, 2019], Chu et al. [2020]
@@
Table 1: Past studies on transport of colloids into and out of the dead-end pore driven by diffusiophoresis and diffusioosmosis. Here, $\zeta_{p}$ (mV), $\zeta_{w}$ (mV), $c^{\ast}$ (mM), and $\beta$ are the particle zeta potential, wall zeta potential, solute concentration in the pore, and solute concentration ratio between reservoir and pore, respectively. (Exp. - experimentally, Num. - numerically). Here Pe is the solute Peclet number, DP and DO represent diffusiophoresis and diffusioomsosis, respectively.
@@
Most of these applications involve transport of colloids within stagnant dead-end pores that facilitate negligible fluid convection. Hence, electrolytic DP can serve as an efficient method for microporous transport. To set this up as colloidal injection or withdrawal, the chemical gradients can be applied in either ‘solute-out’ mode (solute-saturated pore meets solute-free ambient) or ‘solute-in’ mode (solute-free pore meets solute-saturated ambient). In their seminal experimental study, Kar et al. [2015] and Shin et al. [2016] demonstrated how droplets and particles can be efficiently transported in dead-end pores in a controlled manner. Using a continuum framework, Ault et al. [2017] later focused on modeling the colloidal injection and withdrawal dynamics and showed that colloids enter the pores as propagating bands. In the limit of times smaller than characteristic solute diffusion, zero colloidal diffusivity, and weak phoretic mobility, they found that injection wavefronts develop as $\sim t^{1/2}$ for solute-out mode, whereas solute-in mode dynamics show weak temporal development. Later, Gupta et al. [2020] showed the impact of solute concentration on mobility and solute-out compaction, emphasizing the importance of finite-Debye layer effects for transport of sub-micron sized colloids. As part of their study, Lee et al. [2023] further extended the above two studies (for solute-out injection) to large zeta potentials, such that the colloid’s phoretic mobility varies with local electrolyte concentration, i.e., the constant charge (or variable mobility) model. Of the two effects, finite-Debye layer effects have a greater impact on the quantitative nature of the results, whereas their influence on the qualitative nature of the injection profile is negligible.
@@
In addition to particle diffusiophoresis, bulk osmotic flow can shape the particle dispersion because channel or pore walls are generally charged and, in the presence of external solute gradients, can induce diffusio-osmotic slip velocity Kar et al. [2015], Shin et al. [2016]. This slip velocity drives a nonuniform flow within the pores to maintain zero net flow. Using Taylor dispersion analysis and experiments, Alessio et al. [2022] further demonstrated that osmotic flows originating from side walls of the channel can have a significant impact on colloidal dispersion. Recently, a similar experimental and numerical study by Akdeniz et al. [2023] highlighted the impact of implementing constant-charge model on colloidal dispersion, particularly over long timescales. Table 1 provides an overview of these significant contributions towards understanding colloidal transport in dead-end pore.
@@
Since the first experiments and computations Kar et al. [2015], Shin et al. [2016], Ault et al. [2017] on dead-end pores, significant progress has been made in understanding the impact of multiple and multivalent electrolyte gradients Alessio et al. [2021], Gupta et al. [2019], pH gradients Shim et al. [2022], photocatalytic particles Visan and Lammertink [2019], temporal control of solutal conditions Ha et al. [2019], Migacz et al. [2024], and polymeric gradients Akdeniz et al. [2024]. However, these studies have focused primarily on colloidal injection and compaction in the solute-out mode, as they provide the most pronounced experimental results in short timescales, leaving the other three cases largely unexplored. Furthermore, phoretic transport in naturally occurring scenarios Lager et al. [2008], Ganguly et al. [2023], Ault and Shin [2024] may not be limited to the solute-out mode. In this work, we examine previously overlooked qualitative differences and demonstrate that solute-in and solute-out modes of colloidal transport exhibit distinct spatiotemporal dynamics as well as performance characteristics; a preliminary insight can be found in Fig. 1. Specifically, this study is driven by the following questions:
@@
In what follows, we describe the continuum framework used to predict and obtain qualitative insights into colloidal transport. Since the two modes are found to evolve on different time scales, we realize that their lifespan cannot be characterized by the solute-diffusion timescale $t_{d}$ and effectiveness cannot be captured by phoretic velocity $\bm{u}_{p}$, and thus quantify the performance in two metrics: persistence time and effectiveness. Subsequently, we explore the influence of variable mobility model on spatiotemporal patterns. Additionally, we investigate how osmotic flows affect injection and withdrawal dispersion, finding that osmosis-induced mixing can counterintuitively inhibit injection effectiveness in solute-out mode. Finally, we discuss key conclusions and outlook.
II Model Description
We consider the diffusiophoretic motion of freely suspended and non-interacting colloids in dead-end pores driven by the two orientational modes of solute gradients. We consider gradients of symmetric electrolytes, where the ionic concentration is represented by $c(x,t)$. Fig. 1 (a) depicts the case when colloids are withdrawn from a pore, whereas Fig. 1 (b) shows the injection. At the continuum scale, this coupled transport is governed by the following convection-diffusion equations for solutal and colloidal dynamics Shin et al. [2016]
@@ II Model Description
These equations have been non-dimensionalized using $c_{\text{i}}$, $n_{\text{i}}$, $L$, $L^{2}/D_{\text{s}}$, $\Gamma_{w}/L$, and $\mu\Gamma_{w}/L^{2}$ for solute concentration, particle concentration, length, time, velocity, and pressure, respectively. The normalized phoretic velocity $\bm{u}_{p}=\frac{\Gamma_{p}}{D_{\text{s}}}\bm{\nabla}\ln c$ couples the two equations. These equations are complemented with the quasi-steady flow Stokes equations that govern the 2D osmotic flow emerging from pore walls: $\bm{\nabla}\cdot\bm{u_{\text{f}}}=0$ and $0=\bm{\nabla}p+{\bm{\nabla}}^{2}\bm{u_{\text{f}}}$. The solute Peclet number $\text{Pe}=\Gamma_{w}/D_{s}$ captures the ratio of osmotic flow convection relative to solute diffusion. We have defined the dimensionless $\mathcal{D}$ for capturing colloidal diffusion ($D_{\text{p}}/D_{\text{s}}$) relative to the solute diffusion. Here, $c_{\text{i}}$ represents the initial solute concentration in the pore (reservoir) for solute-out (solute-in) mode, and similarly, $n_{\text{i}}$ represents the initial colloidal concentration in the pore (reservoir) for withdrawal (injection).
@@
Figure 2: Spatio-temporal profile of colloids. (a) Pure diffusion profile ($\Gamma_{p}=0$) for injection (solid) and withdrawal (dash-dotted). (b) Dynamics of colloid withdrawal, where solid lines depict shallow withdrawal in solute-out mode and dashed lines depict deeper withdrawal in solute-in mode. (c) Weighted average position of withdrawn population for $\beta=0.05,0.1,0.2,5,10,20$. The dash-dotted line represents bare colloid diffusion for $\beta=1$. (d) Injection dynamics. The arrows depict the axis each mode corresponds to. Parameters: $D_{s}=10^{-9}$ m2/s, $D_{p}=10^{-12}$ m2/s, $\Gamma_{p}=10^{-9}$ m2/s, $\beta=0.1$ (solute-out, solid lines), and $\beta=10$ (solute-in, dashed lines).
@@
Solute concentration can be derived analytically using separation of variables Shin et al. [2016], Ault et al. [2017]
$c(x,t)=\beta+(1-\beta)\sum_{n=0}^{\infty}\frac{2(1-\cos\lambda_{n})}{\lambda_{n}}\sin\left({\lambda_{n}x}\right)e^{-\lambda_{n}^{2}t}.$
@@
We use the first 500 terms of this series to evaluate the phoretic velocity. The 1D colloidal concentration profile was obtained numerically by employing the Method of lines, that utilized second order finit-difference discretization in space. In addition to the grid independence test, we verified our results using COMSOL multiphysics (see §1 in Supplementary material), which is also used to obtain the 2D profile.
III Results
We begin the discussion with the assumption that colloids and pore walls exhibit uniform and constant mobilities. Although zeta potentials can be sensitive to local concentrations Kirby and Hasselbrink Jr [2004], we first focus on building a systemic understanding and relax this assumption in a later subsection. Furthermore, to isolate the phoretic effects for clarity, we consider zero mobility of the confining pore walls, i.e., absence of diffusioosmotic flow ($\text{Pe}\to 0$), which renders the system primarily 1D transient. This assumption will also be relaxed in the subsequent subsections.
III.A Phoretic withdrawal and injection for constant mobility
@@
Withdrawal. First, we explore the withdrawal of colloids that are trapped within the pore at time $t=0$, as depicted in Fig.1(a). Fig.2(a) shows the spatio-temporal colloid density profiles for pure colloidal diffusion ($\Gamma_{p}=0$). It requires $10^{2}$ orders of solute diffusion time ($t_{d}$) to withdraw a substantial amount of particles. Fig. 2(b) shows that diffusiophoresis can offer similar magnitudes of withdrawal in $\sim t_{d}$ times. Estimations of such an intensified transport in dead-end pores have been explored earlier Ault et al. [2017], Lee et al. [2023], Gupta et al. [2020], Ault and Shin [2024]. Here, we focus on the influence electrolytic-gradient direction has on the qualitative and quantitative aspects. The trends in Fig.2(b) depict that solute-out mode offers a shallow withdrawal where colloids from pore-end experience delayed entrainment, whereas solute-in mode facilitates a deeper withdrawal. We quantify these trends in Fig. 2(c) which shows the weighted average position of the withdrawn colloids: $\langle x\rangle_{\text{with}}=1-\left({\int_{0}^{1}n\,x\,\text{d}x}/{\int_{0}^{1}n\,\text{d}x}\right)$, where the measure for solute-in (-out) mode lies above (below) $x=0.5$. Post diffusiophoretic lifespan (after few $t_{d}$), both trends asymptotically approach the diffusion-mediated colloid withdrawal shown by the dash-dotted curve for $\beta=1$.
@@
Figure 4: Performance measures. (a) Variation of phoretic persistence time with applied concentration ratio $\beta$. The dashed lines show the analytical predictions from Eq. (3); $\lambda=\pi^{2}/4$. Persistence is identical for colloid injection and withdrawal. (b) Variation of the diffusiophoretic velocity (at $x=0.5$) with time for varying $\beta$ (for solute-out mode). (c,d) Temporal evolution of effectiveness for withdrawal (c) and injection (d). Solute-in (dashed line) and solute-out (solid line) modes are shown for various magnitudes of applied solute gradients at $D_{s}=10^{-9}$ m2/s, $D_{p}=10^{-12}$ m2/s, and $\Gamma_{p}=10^{-9}$ m2/s.
@@
III.A.1 Persistence time
Given these differences between the two modes, it becomes evident that the lifespan of phoretic transport cannot be adequately characterized by $t_{d}$. Hence, we define persistence time that is calculated as the time taken for $\bm{u}_{p}$ to decay to 99%, averaged over the entire pore length: $\langle\tau\rangle=\frac{1}{L}\int_{0}^{1}\tau(x),\,\text{where }\tau$ the local persistence is defined such that $\bm{u}_{p}(x,\tau(x))=0.01\,\bm{u}_{p}(x,0)$. This gives us an average measure of lifetime of diffusiophoretic process and is identical for injection and withdrawal. Figure 4 (a) shows its variation with the applied concentration difference, a ratio characterized by $\beta$ whose variation spans for both solute-out ($\beta<1$) and solute-in modes ($\beta>1$). We note that solute-out diffusiophoretic modes are relatively more persistent and are quite sensitive to the externally applied gradient: $\tau$ rises logarithmically with diminishing $\beta$. This can be directly deduced by simplifying the phoretic velocity for times $t\gtrsim t_{d}$, enabling a one-term approximation in the infinite series solution for the solute profile, which yields:
$\bm{u}_{p}=\frac{\Gamma_{p}\cos(\pi x/{2})}{e^{\lambda t}(\frac{\beta}{1-\beta})+\frac{4}{\pi}\sin(\frac{\pi x}{2})},$

### v2

@@ Abstract
Abstract. Electrolytic diffusiophoresis refers to directional migration of colloids due to interfacial forces that develop in response to local electrolytic concentration ($c$) gradients. This physicochemical transport provides an efficient alternative in numerous microscale applications where advection-induced transport is infeasible. Phoretic withdrawal and injection in dead-end pores can be controlled by orienting salt gradients into or out of the pore; however, the extent to which this orientation influences spatiotemporal transport patterns is not thoroughly explored. In this study, we find that it has a significant influence: colloidal withdrawal in solute-out mode ($\beta=c_{\infty}/c_{\text{pore}}<1$) is faster and shallower, whereas the solute-in mode enables deeper withdrawal. Similarly, solute-out injection features rapidly propagating wavefronts, whereas the solute-in mode ($\beta>1$) promotes uniform and gradual injection. Each mode’s transport is found to evolve and persist over different time scales. We characterize the performance of these modes and find that while persistence of the solute-out mode strengthens with a growing electrolytic gradient [$\sim\ln(\beta^{-0.4})$], solute-in mode diminishes and eventually its persistence is insensitive to $\beta$. We also incorporate the variable mobility model to examine the impact of large zeta potentials, which intensifies the transport of solute-out mode further and weakens the solute-in mode. Additionally, we investigate how osmotic flows of the two modes affect injection and withdrawal patterns. We find that osmosis-induced mixing can counterintuitively inhibit injection effectiveness in solute-out mode. These insights bring attention to the distinctions between different phoretic transport modes and contribute to the rational design and setup of electrolytic gradients in numerous microscale applications.
Introduction
@@ 0.2-1 $\mu$m polystyrene and
Reassessed refs.10, 3, 11, 12
@@
Table 1: Past studies on transport of colloids into and out of the dead-end pore driven by diffusiophoresis and diffusioosmosis. Here, $\zeta_{p}$ (mV), $\zeta_{w}$ (mV), $c^{\ast}$ (mM), and $\beta$ are the particle zeta potential, wall zeta potential, solute concentration in the pore, and solute concentration ratio between reservoir and pore, respectively. (Exp. - experimentally, Num. - numerically). Here Pe is the solute Peclet number, DP and DO represent diffusiophoresis and diffusioomsosis, respectively. Here $L$ denotes the pore length.
@@
Microscopic colloids in a molecular bath experience potential forces from intermolecular interactions within the Debye screening region near their surface. External concentration gradients disrupt the equilibrium of thermal and interactive forces, inducing a diffusioosmotic (DO) slip velocity. This slip on a freely suspended colloid drives a diffusiophoretic (DP) motion, where it migrates parallel to chemical gradients $\propto\nabla c$ 14, 15. When the solute is an electrolyte, differential ion diffusion creates electrostatic gradients ($\nabla\psi$) alongside chemical gradients. This generates a local electric field that modifies the osmotic/phoretic contribution, proportional to $\nabla\log c$ 16, 17, 18. This motion is mediated by diffusive and electrostatic solute-surface interactions, encapsulated in chemo- and electro-phoretic mobility parameters.
The resurgence of these phoretic and osmotic phenomena after decades of initial work is attributed to the rise of micro- and nano-fluidic technologies, which enable precise control over fluid and particle transport 19. Through experiments on focusing and spreading of colloids, the seminal study by Abécassis et al. 20 demonstrated two orders-of-magnitude enhancement in effective ‘diffusion’ in salt gradients $\sim\Gamma_{p}^{2}/D_{s}$ ($\Gamma_{p}$ and $D_{s}$ are phoretic mobility and salt diffusion coefficient). Later, Palacci et al. 21 built microfluidic devices for patterned trapping of colloids and $\lambda-$DNA in temporally-controlled soute concentration profile. In response to salt oscillations at varying frequencies, the particles settled into a characteristic Gaussian distribution. The width of this distribution depended on the phoretic mobility, which varied with different salts (LiCl > NaCl > KCl), as it is influenced by the difference in the diffusivities of the dissociated ions. Later, Palacci et al. 22 analysed the time scales and robust persistence of colloidal migration via experiments and numerical simulations. They demonstrated that phoretic log-sensing enables efficient particle focusing across various “osmotic shock” configurations within a microfluidic circular well chamber. These physicochemical effects can also be leveraged for designing bioassays that trap vesicles 23 & transport DNA 24, performing low-cost zeta potentiometry 4, 11, and designing hydrophobic nanopores that can act as osmotic diodes for rectified transport 25. Furthermore, potential discoveries of osmotic and phoretic phenomena have recently emerged within applications and technologies that have existed for decades. Examples include, dirt transport in constricted fibre pores during conventional laundering 26, phoresis-induced fouling, osmotic flows in water filtration membranes 27, 28, and low-salinity water flooding for tertiary oil recovery 1, 17, 29. Lee et al. 30 fabricated a microfluidic device for continuous water purification in conduits made from ion-exchange membranes and studied exclusion of particle and pathogens in a developing ionic boundary layer. Bekir et al. 31 employed light-induced diffusioosmosis on particle surfaces to yield a lift, which aided in separating particles on the basis of size and porosity. Chakra et al. 32 devised a continuous experimental procedure to sort and characterize nanoparticles and liposomes on a $\Psi$-junction microfluidic chip.
@@
Most of these applications involve transport of colloids within stagnant dead-end pores that facilitate negligible fluid convection. Hence, electrolytic DP can serve as an efficient method for microporous transport. To set this up as colloidal injection or withdrawal, the chemical gradients can be applied in either ‘solute-out’ mode (solute-saturated pore meets solute-free ambient) or ‘solute-in’ mode (solute-free pore meets solute-saturated ambient). In their seminal experimental study, Kar et al. 1 and Shin et al. 2 demonstrated how droplets and particles can be efficiently transported in dead-end pores in a controlled manner. Using a continuum framework, Ault et al. 3 later focused on modeling the colloidal injection and withdrawal dynamics and showed that colloids enter the pores as propagating bands. In the limit of times smaller than characteristic solute diffusion, zero colloidal diffusivity, and weak phoretic mobility, they found that injection wavefronts develop as $\sim t^{1/2}$ for solute-out mode, whereas solute-in mode dynamics show weak temporal development. Later, Gupta et al. 6 showed the impact of solute concentration on mobility and solute-out compaction, emphasizing the importance of finite-Debye layer effects for transport of sub-micron sized colloids. As part of their study, Lee et al. 9 further extended the above two studies (for solute-out injection) to large zeta potentials, such that the colloid’s phoretic mobility varies with local electrolyte concentration, i.e., the constant charge (or variable mobility) model. Of the two effects, finite-Debye layer effects have a greater impact on the quantitative nature of the results, whereas their influence on the qualitative nature of the injection profile is negligible. Recently, Tan et al. 33 demonstrated a two-step process for persistent delivery of colloids in porous media by utilizing targeted dead-end pores—i.e., pore ends laden with high-solute-absorption-capacity, such as Polyethylene Glycol Diacrylate (PEG-DA). In the first step, the entire pore (including the target) are saturated with solute. In the second step, a solute-free colloidal suspension encounters the pore mouth. Due to the high partition coefficient of PEG-DA target, it acts as a beacon, slowly releasing solute and maintaining long-lasting gradients that facilitate deeper colloid injection.
@@
In addition to particle diffusiophoresis, bulk osmotic flow can shape the particle dispersion because channel or pore walls are generally charged and, in the presence of external solute gradients, can induce diffusio-osmotic slip velocity 1, 2. This slip velocity drives a nonuniform flow within the pores to maintain zero net flow. Using Taylor dispersion analysis and experiments, Alessio et al. 7 further demonstrated that osmotic flows originating from side walls of the channel can have a significant impact on colloidal dispersion. Recently, a similar experimental and numerical study by Akdeniz et al. 8 highlighted the impact of implementing constant-charge model on colloidal dispersion, particularly over long timescales. Table 1 provides an overview of these significant contributions towards understanding colloidal transport in dead-end pore.
@@
Since the first experiments and computations 1, 2, 3 on dead-end pores, significant progress has been made in understanding the impact of multiple and multivalent electrolyte gradients 34, 35, pH gradients 36, photocatalytic particles 37, temporal control of solutal conditions 38, 13, and polymeric gradients 39. However, these studies have focused primarily on colloidal injection and compaction in the solute-out mode, as they provide the most pronounced experimental results in short timescales, leaving the other three cases largely unexplored. Furthermore, phoretic transport in naturally occurring scenarios 29, 40, 18 may not be limited to the solute-out mode. In this work, we examine previously overlooked qualitative differences and demonstrate that solute-in and solute-out modes of colloidal transport exhibit distinct spatiotemporal dynamics as well as performance characteristics; a preliminary insight can be found in Fig. 1. Specifically, this study is driven by the following questions:
@@
In what follows, we describe the continuum framework used to predict and obtain qualitative insights into colloidal transport. Since the two modes evolve over vastly different timescales, their lifespans cannot be characterized by a common phoretic timescale $L^{2}/\Gamma_{p}$. We therefore quantify performance using two separate metrics: persistence time and effectiveness. Subsequently, we explore the influence of variable mobility model on spatiotemporal patterns. Additionally, we investigate how osmotic flows affect injection and withdrawal dispersion, finding that osmosis-induced mixing can counterintuitively inhibit injection effectiveness in solute-out mode. Finally, we discuss key conclusions and outlook.
Materials and Methods
Mathematical Model. We consider the diffusiophoretic motion of freely suspended and non-interacting colloids in dead-end pores driven by the two orientational modes of solute gradients. We consider gradients of symmetric electrolytes, where the ionic concentration is represented by $c(x,t)$. Fig. 1 (a) depicts the case when colloids are withdrawn from a pore, whereas Fig. 1 (b) shows the injection. At the continuum scale, this coupled transport is governed by the following convection-diffusion equations for solutal and colloidal dynamics 2
@@
These equations have been non-dimensionalized using $c_{\text{i}}$, $n_{\text{i}}$, $L$, $L^{2}/D_{\text{s}}$, $\Gamma_{w}/L$, and $\mu\Gamma_{w}/L^{2}$ for solute concentration, particle concentration, pore length, time, velocity, and pressure, respectively. The normalized phoretic velocity $\bm{u}_{p}=\frac{\Gamma_{p}}{D_{\text{s}}}\bm{\nabla}\ln c$ couples the two equations. These equations are complemented with the quasi-steady flow Stokes equations that govern the 2D osmotic flow emerging from pore walls: $\bm{\nabla}\cdot\bm{u_{\text{f}}}=0$ and $0=\bm{\nabla}p+{\bm{\nabla}}^{2}\bm{u_{\text{f}}}$. The solute Peclet number $\text{Pe}=\Gamma_{w}/D_{s}$ captures the ratio of osmotic flow convection relative to solute diffusion. We have defined the dimensionless $\mathcal{D}$ for capturing colloidal diffusion ($D_{\text{p}}/D_{\text{s}}$) relative to the solute diffusion. Here, $c_{\text{i}}$ represents the initial solute concentration in the pore (reservoir) for solute-out (solute-in) mode, and similarly, $n_{\text{i}}$ represents the initial colloidal concentration in the pore (reservoir) for withdrawal (injection).
@@
Figure 2: Spatio-temporal profile of colloids. (a) Pure diffusion profile ($\Gamma_{p}=0$) for injection (dash-dotted) and withdrawal (solid). (b) Dynamics of colloid withdrawal, where solid lines depict shallow withdrawal in solute-out mode and dashed lines depict deeper withdrawal in solute-in mode. (c) Weighted average position of withdrawn population for $\beta=0.05,0.1,0.2,5,10,20$. Here, solid and dashed lines depict the solute-out and solute-in modes, respectively. The dash-dotted line represents bare colloid diffusion for $\beta=1$. (d) Injection dynamics. The arrows depict the axis each mode corresponds to. Parameters: $D_{s}=10^{-9}$ m2/s, $D_{p}=10^{-12}$ m2/s, $\Gamma_{p}=10^{-9}$ m2/s, $\beta=0.1$ (solute-out, solid lines), and $\beta=10$ (solute-in, dashed lines).
@@
Computational Method. Solute concentration can be derived analytically using separation of variables 2, 3
$c(x,t)=\beta+(1-\beta)\sum_{n=0}^{\infty}\frac{2(1-\cos\lambda_{n})}{\lambda_{n}}\sin\left({\lambda_{n}x}\right)e^{-\lambda_{n}^{2}t},$ (3)
@@
where $\lambda_{n}=\pi(2n+1)/2$. We use the first 103 terms of this series to evaluate the phoretic velocity. We note that the step change in solute concentration at pore entry corresponds to a singularity in phoretic velocity at $t=0$. For practical purposes, we focus on the times $t^{*}\geq 0.1t_{d}$ in all our numerical results. In employing the series solution for solute, we use the first 103 terms to ensure excellent accuracy. The 1D colloidal concentration profile was obtained numerically by employing the Method of lines, that utilized second order finite-difference discretization in space ($\Delta x=0.001$ $\Delta t=0.005$). In addition to the grid independence test, we verified our results using COMSOL multiphysics (see §1 in Supplementary material), which is also used to obtain the 2D profile. Therein, we also compare our results with earlier numerical studies 2, 9.
Results and Discussion
We begin the discussion with the assumption that colloids and pore walls exhibit uniform and constant mobilities. Although zeta potentials can be sensitive to local concentrations 41, we first focus on building a systemic understanding and relax this assumption in a later subsection. Furthermore, to isolate the phoretic effects for clarity, we consider zero mobility of the confining pore walls, i.e., absence of diffusioosmotic flow ($\text{Pe}\to 0$), which renders the system primarily 1D transient. This assumption will also be relaxed in the subsequent subsections. In the results below, we consider colloids with both positive and negative phoretic mobilities, which can arise for either positive or negative zeta potentials, provided the ion diffusivity difference, defined as $\beta_{s}=\frac{D_{+}-D_{-}}{D_{+}+D_{-}}$, has an appropriate sign and magnitude. Using the mobility expression from Prieve et al. 16 and assuming room temperature properties of water, estimates from the experimental studies of Shin et al. 4 suggest that a micron-sized native polystyrene particle ($\zeta_{p}=-79\,\text{mV}$) yields a mobility of $\Gamma_{p}=0.95\times 10^{-9}\,\text{m}^{2}/\text{s}$ in NaCl ($\beta_{s}=-0.208$), whereas in SDS ($\beta_{s}=0.55$), the mobility is approximately $-0.26\times 10^{-9}\,\text{m}^{2}/\text{s}$. Although positive mobility is more common in neutral electrolytes like NaCl, polystyrene particles in concentrated HCl mediums ($\zeta_{p}=-34$ mV, $\beta_{s}=0.642$) can yield a negative mobility of $-0.33*10^{-9}$ m2/s 36. Furthermore, amine-modified polystyrene particles can exhibit positive zeta potentials ($\zeta_{p}\approx+43$ mV) that correspond to negative mobilities of $-0.34*10^{-9}$ m2/s in NaOH gradients ($\beta_{s}=-0.596$) 36. Following a similar rationale, we also consider both positive and negative mobilities for pore walls in a later subsection. Although direct experimental evidence for positive wall zeta potentials or negative wall mobilities is currently lacking, such conditions are theoretically plausible based on the particle-level arguments. We note that cases involving particles and walls with opposite zeta potentials are experimentally challenging due to electrostatic-interaction-induced fouling 42.
Phoretic withdrawal and injection for constant mobility
@@
Withdrawal. First, we explore the withdrawal of colloids that are trapped within the pore at time $t=0$, as depicted in Fig.1(a). Fig.2(a) shows the spatio-temporal colloid density profiles for pure colloidal diffusion ($\Gamma_{p}=0$). It requires $10^{2}$ orders of solute diffusion time ($t_{d}$) to withdraw a substantial amount of particles. Fig. 2(b) shows that diffusiophoresis can offer similar magnitudes of withdrawal in $\sim t_{d}$ times. Estimations of such an intensified transport in dead-end pores have been explored earlier 21, 22, 9, 6, 18. Here, we focus on the influence electrolytic-gradient direction has on the qualitative and quantitative aspects. The trends in Fig.2(b) depict that solute-out mode offers a shallow withdrawal where colloids from pore-end experience delayed entrainment, whereas solute-in mode facilitates a deeper withdrawal. We quantify these trends in Fig. 2(c) which shows the weighted average position of the withdrawn colloids: $\langle x\rangle_{\text{with}}=1-\left({\int_{0}^{1}n\,x\,\text{d}x}/{\int_{0}^{1}n\,\text{d}x}\right)$, where the measure for solute-in (-out) mode lies above (below) $x=0.5$. Post diffusiophoretic lifespan (after few $t_{d}$), both trends asymptotically approach the diffusion-mediated colloid withdrawal shown by the dash-dotted curve for $\beta=1$.
@@
Figure 4: Performance measures. (a) Variation of phoretic persistence time with applied concentration ratio $\beta$. The purple lines at the extremes show the analytical predictions from Eq. (5); $\lambda=\pi^{2}/4$. Persistence is identical for colloid injection and withdrawal. The empty circles correspond to Eq. (6). (b) Variation of the diffusiophoretic velocity (at $x=0.5$) with time for varying $\beta$ (for solute-out mode). (c,d) Temporal evolution of effectiveness for withdrawal (c) and injection (d). Solute-in (dashed line) and solute-out (solid line) modes are shown for various magnitudes of applied solute gradients at $D_{s}=10^{-9}$ m2/s, $D_{p}=10^{-12}$ m2/s, and $\Gamma_{p}=10^{-9}$ m2/s. Here, (c) and (d) share the same legends for $\beta$.
@@
Persistence time
Given these differences between the two modes, it becomes evident that the lifespan of phoretic transport cannot be adequately characterized by $t_{d}$. Hence, we define persistence time that is calculated as the time taken for $\bm{u}_{p}$ to decay to 99% from its maximum value, averaged over the entire pore length: $\langle\tau\rangle=\frac{1}{L}\int_{0}^{1}\tau(x),\,\text{where }\tau$ the local persistence is defined such that $\bm{u}_{p}(x,\tau(x))=0.01\,\bm{u}_{p}(x,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.1})$. In our calculations, we consider earliest time to be $t^{*}=0.1t_{d}$. This definition gives us an average measure of lifetime of diffusiophoretic process and is identical for injection and withdrawal. Figure 4 (a) shows its variation with the applied concentration difference, a ratio characterized by $\beta$ whose variation spans for both solute-out ($\beta<1$) and solute-in modes ($\beta>1$). We note that solute-out diffusiophoretic modes are relatively more persistent and are quite sensitive to the externally applied gradient: $\tau$ rises logarithmically with diminishing $\beta$. This can be directly deduced by simplifying the phoretic velocity for times $t\gtrsim t_{d}$, enabling a one-term approximation in the infinite series solution for the solute profile, which yields:
$\bm{u}_{p}=\frac{\Gamma_{p}\cos(\pi x/{2})}{e^{\lambda t}(\frac{\beta}{1-\beta})+\frac{4}{\pi}\sin(\frac{\pi x}{2})},$ (4)